OpenStudy (ytry):

A balloon is a sphere with a radius of 5.0 m. The force of air against the walls of the balloon is 45 N. What is the air pressure inside the balloon? (surface area = 12.56 x radius2; 1 Pa = 1 N/m2) This is my question for my homework I am unsure how to solve it.

mathslover (mathslover):

Pressure = Force / Area Given Force = 45 N Radius = 5 m Therefore Surface area becomes : 12.56 * (Radius)^2 = 12.56 * (5 m)^2 You have area and force now, calculate the pressure : \(Pressure = \cfrac{45 N}{12.56 \times (5 m)^2}\)

OpenStudy (anonymous):

Yeah mathslover is right. 45N/314m^2 = .143 Pa

mathslover (mathslover):

Yep. ^